# A parametric study of the pipeline hammer phenomenon in plastic Bingham slurry flows using the finite element method

### v1

@@ 2.3 Dimensionless equations
The system given in Eq.(2.1) can be expressed in dimensionless form by introducing relevant characteristic scales for the velocity ($U_{o}$), pressure ($P_{o}$), position ($L_{o}$) and time ($T_{o}$) [29, 30, 24]. The scales for velocity and pressure were already introduced in the system (2.14). Exploiting the simplicity of the experimental setup, the time scale is chosen as $T_{o}=L/c$, for $L_{o}=L$. These definitions leads to the unknowns $x^{*},t^{*},u^{*},p^{*}$ defined as:
@@
$\frac{\partial p^{*}}{\partial t^{*}}+\lambda_{1}\frac{\partial u^{*}}{\partial x}&=0&\text{ in }(x^{*},t^{*})\in[0,1]\times[0,T^{*}],\\ \lambda_{1}\frac{\partial u^{*}}{\partial t^{*}}+\lambda_{2}u^{*}\frac{\partial u^{*}}{\partial x^{*}}+\frac{\partial p^{*}}{\partial x^{*}}+\frac{f\text{Re}}{16}u^{*}&=0&\text{ in }(x^{*},t^{*})\in[0,1]\times[0,T^{*}],\\ u^{*}&=g^{*}(t)&\text{ on }x^{*}=1,~{}t^{*}\geq 0,\\ p^{*}&=1&\text{ on }x=0,~{}t^{*}\geq 0,\\ p^{*}&=1-x^{*}&\text{ at }t^{*}=0,~{}x^{*}\in[0,1],\\ u^{*}&=\frac{16}{f\widetilde{\text{Re}}}&\text{ at }t^{*}=0,~{}x^{*}\in[0,1],\end{aligned}\right.$ (2.14)
@@
$\lambda_{1}$ $=\frac{\delta}{32}\widetilde{\text{Re}}~{}\widetilde{\text{Ma}}^{-1}$ (2.15a)
$\lambda_{2}$ $=\frac{\delta}{32}\widetilde{\text{Re}}$ (2.15b)
@@ 2.4 A lowest-order finite element formulation for slurry flows
Traditionally, 1D hammer dynamics have been solved by means of the Method of Characteristics (MOC). The MOC transforms the partial differential equations (PDEs) into ordinary differential equations (ODEs) along characteristic lines, which represent the paths along which wave information travels through the fluid. This method allows for the accurate simulation of wave propagation, reflection, and interaction within the pipeline system. Nevertheless, MOC relies upon restrictive hypothesis for the underlying model that force a set of boundaries for the applications to be tackled. For a system of continuous equations to be addressed with MOC, it is mandatory to satisfy quasi-linear conditions, meaning that in the equations one can allow either linear terms or, at most, terms that depends solely on the unknowns magnitude and not their derivatives. This is of relevance, for instance, whenever it is desired to solve the system considering complex friction models.
Among the alternatives to perform to numerically solve in space and time the equations (2.14) we can name the MOC [24], finite volumes [43], and the finite element method (FEM, [44] [45]). We will use the later to deal with space discretization and implicit finite differences for time. Let us consider a one-dimensional working domain, a straight line represented by the real interval $\Omega=[0,L]$ ($L\in\mathbb{R}$). In order to properly expose the numerical setup, it is required to formalize the functions $u^{*}:[0,L]\rightarrow\mathbb{R}$ and $p^{*}:[0,L]\rightarrow\mathbb{R}$ as an element of a Hilbert space $\mathcal{H}=H^{1}([0,L])\times L^{2}([0,L])$, where $L^{2}=\{v:\Omega\rightarrow\Omega:\int_{\Omega}\lvert v\rvert^{2}\mathrm{d}x<\infty\}$, and $H^{1}(\Omega)$ stands for the smaller space of smoother functions $H^{1}=\{v\in L^{2}(\Omega);~{}\partial v/\partial x\in L^{2}(\Omega)\}$.
Under this setting, it is well known that one can reformulate the problem with a variational formulation, boiling the challenge down to finding a set of functions $(u^{*,n+1},p^{*,n+1})\in\mathcal{H}$, such that:
$\frac{1}{\Delta t}\int_{\Omega}p^{*,n+1}q\mathrm{d}x+\lambda_{1}\int_{\Omega}\frac{\partial u^{*,n+1}}{\partial x^{*}}q\mathrm{d}x+\frac{\lambda_{1}}{\Delta t}\int_{\Omega}u^{*,n+1}w\mathrm{d}x+\int_{\Omega}\frac{\partial p^{*,n+1}}{\partial x^{*}}w+$ (2.16)
$+\lambda_{2}\int_{\Omega}u^{*}\frac{\partial u^{*}}{\partial x^{*}}w\mathrm{d}x+\frac{f\text{Re}}{16}\int_{\Omega}u^{*,n+1}w\mathrm{d}x=\frac{1}{\Delta t^{*}}\int_{\Omega}p^{*,n}q\mathrm{d}x+\frac{\lambda_{1}}{\Delta t^{*}}\int_{\Omega}u^{*,n}w\mathrm{d}x,$
@@
which means that we need to solve the system (2.17) consecutively to complete the simulation time with a uniform grid $[0,\Delta{t},2\Delta{t},\ldots]$. A traditional Picard method is used to deal with the system non-linearities at each time-step, which will have a fixed error tolerance over the $\ell^{2}$ norm for joint velocity and pressure of $10^{-3}$. Next, we recast the problem in finite dimension by means of the following standard Galerkin approach:
@@
where $N$ is the number of nodes that discretize the working domain $\Omega$. To ease the reading, the same notation has been adopted for both the continuous and the discrete unknowns. The finite element method provides the solution of nodal coefficients $u^{*,n+1}_{i}$ and $p^{*,n+1}$ ($i=1,\ldots,N$). We choose the same discretization space for both, velocity and pressure, namely, the piece-wise linear (lowest-order) space $\mathbb{P}_{1}$ basis $\phi_{1},\ldots,\phi_{N}$. Then, it is easily shown that this approximation leads to the following matrix system of $2N\times 2N$ equations:
$\begin{bmatrix}(1+\Delta t^{*}\frac{f\text{Re}}{16})M+\lambda_{2}\Delta t^{*}C(u^{*,n})\quad&\Delta t^{*}B\\ \lambda_{1}\Delta t^{*}B&M\end{bmatrix}\begin{bmatrix}u^{*,n+1}\\ p^{*,n+1}\end{bmatrix}=\begin{bmatrix}Mu^{*,n}\\ Mp^{*,n}\end{bmatrix}$ (2.19)
@@
where $M\in\mathbb{R}^{N\times N}$ stands for a mass matrix with entries $M_{ij}=\int_{\Omega}\phi_{i}\phi_{j}\mathrm{d}x$, the matrix $B\in\mathbb{R}^{N\times N}$ for the coupling terms, with entries $B_{ij}=\int_{\Omega}\phi_{i}\frac{\partial\phi_{j}}{\partial Z}\mathrm{d}x$, and the convection matrix $C\in\mathbb{R}^{N\times N}$ with entries $C_{ij}=\int_{\Omega}u^{*}\frac{\partial\phi_{i}}{\partial x^{*}}\phi_{j}\mathrm{d}x$. The integrals are computed with a standard Gaussian quadrature rule of three points per finite element. In addition, a classical fixed-point root-finder is utilized to solve for $X_{c}$ in equation (2.12) for a given input Hedstrom number, thus allowing us to compute the yield stress and the adimensional constants $\lambda_{1}$ and $\lambda_{2}$.
@@
The assemblage of the matrices and the solver is handled with the in-house software MAD [46] (used already in several CFD works with finite elements, e.g. [47, 48, 49, 50, 51]), for finite elements and data assimilation, which supports CPU parallelization for optimal speed-up thanks to the underlying linear algebra library PETSc[52]. For this study, we use an AMD Epyc processor with 48 physical cores. We use the MUltifrontal Massively Parallel sparse direct Solver (MUMPs, [53, 54]) to compute the solutions of (2.19), .
@@
The reader may remark the absence of stabilization techniques, as it is the gold standard for the discretization of the Navier-Stokes equations. In particular, we remark on the importance of the mass storage component $(1/\Delta t^{*})M$ in the second line of (2.19) as a regularizer that prevents the system of equations from getting close to singular. One should consider nonetheless that getting close to a divergence-free scenario may distort the condition number of the matrix in (2.19). On the other hand, it would be easy to counter a divergence-free situation. There are several stabilization techniques that would take this situation under control, such as the Brezzi-Pitkäranta approach [55], the streamline upwind Petrov-Galerkin one [56], or variational multi-scale methods [57]. The later will be explored in a fore-coming scientific study to test out unstable regimes ensuring numerical resolvability. In addition, the reader may notice that we follow a monolithic scheme for the problem unknowns. Other decoupling strategies for velocity and pressure for this application may be studied in forthcoming research [58].
@@ 3 Results
This section is devoted to study the numerical solutions of the transient slurry equations with convective effects. We first set up, in section 3.1 the necessary mesh and time step for the finite element scheme we employ so that we reach non-oscillatory or not excessively regularized solutions, as a quasi-convergence study. Next, in section 3.2, we inspect the numerical solution for a realistic parameter set that emulates copper-water mixture. We then do a parametric study of the solutions including the effect of the fluid yield stress (section 3.3) and the impact of the valve closure time (section 3.4).
@@ 3.1 Convergence study
Let start testing out the solver convergence in space and time. To do so, we solve the governing dynamics for the set of parameters depicted in table 1, but considering $\tau_{y}=0$, as it is the more challenging scenario from the numerical standpoint. For similar reasons, we also set a sudden closure for the valve, meaning that in the boundary conditions of (2.14) we set $g(t)=0$ at $x=L$. To study the mesh convergence, we narrow our attention to the pressure at a control point located at the valve. Consider a preliminary adimensional time step of $10^{-6}$ and 4 meshes, with characteristic adimensional element sizes of $10^{-2}$, $2\times 10^{-3}$, $10^{-3}$, and $10^{-4}$. We observe, in Figures 3.a and 3.b, a mesh independent solution up to $\Delta x^{*}=2\times 10^{-3}$, so we choose this element size for all the fore coming simulations. The critical jump is that of the wave onset, at the beginning of the simulation, so we zoom in the phenomena in Figure 3. For every simulation, we check how the reservoir pressure is recovered after the flow stops, as it is physically expected.
Figure 3: Mesh dependence test. As expected, the mesh refinement is critical after the instantaneous valve closure. This sets a threshold for the element size at $\Delta x^{*}=2\times 10^{-3}$
Similarly, the time step analysis in Figure 4.a shows that time steps over $2\times 10^{-3}$ s leads to excessively smooth solutions. This, and the fast Fourier transform of the solutions shown in Figure 4.b depicts a good frequency content preservation and solution behavior for a time step $\Delta t^{*}=10^{-4}$, therefore we use it for all the simulations to come.
@@
Figure 4: Convergence test for Newtonian case $\text{He}=0$. This benchmark shows that $\Delta t=2\times 10^{-5}$ s ($\Delta t^{*}=10^{-4}$) suffices to capture the hammer dynamics.
Concerning the solver performance, which in this test case boils down to invert 50000 systems of equations (one per time step) with 1001 degrees of freedom each, it takes a total simulation time of 66.34 seconds, using a single core of a 48 cores AMD Epyc processor.
@@ 3.2 Simulation of a copper slurry
In this section we will apply the above system of equations to analyze the evolution of a hydraulic transient in a copper-water mixture pipeline system. This mixture is a homogeneous fluid, that is, a fluid whose solid concentration and velocity distribution can be considered practically uniform across the pipe diameter [15, 19]. This case is very important to the mining industry, especially in countries where mineral transport occurs through pipelines.
Table 1: Physical and mechanical parameters for a copper slurry. Viscosity and yield stress data were extracted from [59]
@@
Table 1 shows the physical and mechanical properties for the slurry flow considered in this study, Table 2 the mixture density and bulk modulus and the expected theoretical values for the shockwave speed estimated according to the relations given by Eq.2.8 - Eq.2.10 for increasing values of $C_{V}$. Table 3 display the pipeline properties.
Table 2: Schockwave speeds ($c$) for a copper tailing according to different authors. In the last column, overpressure values ($\Delta{P}$) were included according to Joukowski formula $\Delta P=\rho_{m}c_{m}u_{o}$
@@
Table 3: Parameter setting for the pipeline. The steel pipeline is class DN100, Sch.40s
@@
In agreement with authors experience in real engineering projects, results provided by Eqs. (2.8)-(2.9) are extremely large and then unrealistic. The wave speed values calculated from Eq.(2.10) can be considered most representative of a real transient event, in agreement with Kodura et al. [5, 6, 7]. Even more, it is expected that real values of the initial pressure surge are lower than theoretical estimations inspired on Joukowski’s pressure scale, because of the lower shockwave speed (see e.g., [5, 7]).
@@
With all these elements, the pressure rise can be easily estimated as $\Delta{P}=\rho_{m}c_{m}u_{o}$. The last column of Table 2 shows some estimations of $\Delta P$ by considering $u_{o}=u_{c}$=2.72 m/s a reference velocity. The maximum pressure values $P_{max}$ obtained from numerical simulations can be later compared to the theoretical maximum pressure $P_{o}+\Delta P$, where $P_{o}=\rho_{m}gH$ is the initial head static pressure. As expected, a risky situation for pipeline safety occurs when $P_{max}\geq P_{A}$, where $P_{A}$ is the maximum allowable pressure. This last parameter is provided by the pipeline factory and it depends on wall schedule (see Table 3).
@@
In addition, we show the temporal evolution of the numerical solutions for control points located at the tank ($x=0$), the pipe half length ($x=L/2$), and the valve ($x=L$), at Figures 6 and 7. We observe a rather slow friction dissipation and we check that the maximal numerical pressure reach a value of $13.86$ MPa, which turns out to be surprisingly close to the factory allowable pressure $P_{A}=17.8$ MPa, but still ensuring a safe hammer wave propagation. The theoretical pressure rise discussed above, with more restricted assumptions, is lower than the numerical one, namely $11$ MPa, yet in accordance with the obtained value.

### v2

@@ 2.3 Dimensionless equations
The system given in Eq.(2.1) can be expressed in dimensionless form by introducing relevant characteristic scales for the velocity ($U_{o}$), pressure ($P_{o}$), position ($L_{o}$) and time ($T_{o}$) [31, 32, 26]. The scales for velocity and pressure were already introduced in the system (2.1). Exploiting the simplicity of the experimental setup, the time scale is chosen as $T_{o}=L_{o}/c$, for $L_{o}=L$. These definitions leads to the unknowns $x^{*},t^{*},u^{*},p^{*}$ defined as:
@@
$\frac{\partial p^{*}}{\partial t^{*}}+\lambda_{2}\frac{\partial u^{*}}{\partial x}&=0&\text{ in }(x^{*},t^{*})\in[0,1]\times[0,T^{*}],\\ \lambda_{2}\frac{\partial u^{*}}{\partial t^{*}}+\lambda_{1}u^{*}\frac{\partial u^{*}}{\partial x^{*}}+\frac{\partial p^{*}}{\partial x^{*}}+\frac{f(\text{Re},\text{He})\text{Re}}{16}u^{*}&=0&\text{ in }(x^{*},t^{*})\in[0,1]\times[0,T^{*}],\\ u^{*}&=\frac{16}{f(\widetilde{\text{Re}},\text{He})\widetilde{\text{Re}}}g^{*}(t)&\text{ on }x^{*}=1,~t^{*}\geq 0,\\ p^{*}&=1&\text{ on }x=0,~t^{*}\geq 0,\\ p^{*}&=1-x^{*}&\text{ at }t^{*}=0,~x^{*}\in[0,1],\\ u^{*}&=\frac{16}{f(\widetilde{\text{Re}},\text{He})\widetilde{\text{Re}}}&\text{ at }t^{*}=0,~x^{*}\in[0,1],\end{aligned}\right.$ (2.15)
@@
$\lambda_{1}$ $=\frac{\delta}{32}\widetilde{\text{Re}}$ (2.16a)
$\lambda_{2}$ $=\frac{\delta}{32}\widetilde{\text{Re}}~\widetilde{\text{Ma}}^{-1}$ (2.16b)
@@ 2.4 A lowest-order finite element formulation for slurry flows
Among the alternatives to perform to numerically solve in space and time the equations (2.15) we can name the MOC [26], finite volumes [61], and the finite element method (FEM, [62] [63]). We will use the latter to deal with space discretization and implicit finite differences for time. Let us consider a one-dimensional working domain, a straight line represented by the real interval $\Omega=[0,L]$ ($L\in\mathbb{R}$). In order to properly expose the numerical setup, it is required to formalize the functions $u^{*}:[0,L]\rightarrow\mathbb{R}$ and $p^{*}:[0,L]\rightarrow\mathbb{R}$ as an element of a product Hilbert space $\mathcal{H}=H^{1}([0,L])\times L^{2}([0,L])$, where $L^{2}=\{v:\Omega\rightarrow\Omega:\int_{\Omega}\lvert v\rvert^{2}~\mathrm{d}x<\infty\}$, and $H^{1}(\Omega)$ stands for the smaller space of smoother functions $H^{1}=\{v\in L^{2}(\Omega);~\partial v/\partial x\in L^{2}(\Omega)\}$.
In this setting, the problem can be recast in variational form, reducing the task to finding a pair of functions $(u^{*,n+1},p^{*,n+1})\in\mathcal{H}$ such that:
$\frac{1}{\Delta t^{*}}\int_{\Omega}p^{*,n+1}q~\mathrm{d}x+\lambda_{2}\int_{\Omega}\frac{\partial u^{*,n+1}}{\partial x^{*}}q~\mathrm{d}x+\frac{\lambda_{2}}{\Delta t^{*}}\int_{\Omega}u^{*,n+1}w~\mathrm{d}x+\int_{\Omega}\frac{\partial p^{*,n+1}}{\partial x^{*}}w~\mathrm{d}x+$ (2.17)
$+\lambda_{1}\int_{\Omega}u^{*}\frac{\partial u^{*}}{\partial x^{*}}w~\mathrm{d}x+\frac{f\text{Re}}{16}\int_{\Omega}u^{*,n+1}w~\mathrm{d}x=\frac{1}{\Delta t^{*}}\int_{\Omega}p^{*,n}q~\mathrm{d}x+\frac{\lambda_{2}}{\Delta t^{*}}\int_{\Omega}u^{*,n}w~\mathrm{d}x,$
@@
which means that we need to solve the system (2.18) consecutively to complete the simulation time with a uniform time grid $[0,\Delta{t},2\Delta{t},\ldots]$. A traditional Picard method is used to deal with the system non-linearities at each time-step, which will have a fixed error tolerance over the $\ell^{2}$ norm for joint velocity and pressure of $10^{-3}$. Next, we recast the problem in finite dimension by means of the following standard Galerkin approach:
@@
where $N$ is the number of nodes that discretize the working domain $\Omega$. To ease the reading, the same notation has been adopted for both the continuous and the discrete unknowns. The finite element method provides the solution of nodal coefficients $u^{*,n+1}_{i}$ and $p^{*,n+1}$ ($i=1,\ldots,N$). We choose the same discretization space for both, velocity and pressure, namely, the continuous piece-wise linear (lowest-order) space $\mathbb{P}_{1}$ basis $\phi_{1},\ldots,\phi_{N}$. Then, it is easily shown that this approximation leads to the following matrix system of $2N\times 2N$ equations:
$\begin{bmatrix}(1+\Delta t^{*}\frac{f\text{Re}}{16})M+\lambda_{1}\Delta t^{*}C(u^{*,n})\quad&\Delta t^{*}B\\ \lambda_{2}\Delta t^{*}B&M\end{bmatrix}\begin{bmatrix}u^{*,n+1}\\ p^{*,n+1}\end{bmatrix}=\begin{bmatrix}\lambda_{2}Mu^{*,n}\\ Mp^{*,n}\end{bmatrix}$ (2.20)
@@
where $M\in\mathbb{R}^{N\times N}$ stands for a mass matrix with entries $M_{ij}=\int_{\Omega}\phi_{i}\phi_{j}~\mathrm{d}x$, the matrix $B\in\mathbb{R}^{N\times N}$ for the coupling terms, with entries $B_{ij}=\int_{\Omega}\phi_{i}\frac{\partial\phi_{j}}{\partial Z}~\mathrm{d}x$, and the convection matrix $C\in\mathbb{R}^{N\times N}$ with entries $C_{ij}=\int_{\Omega}u^{*}\frac{\partial\phi_{i}}{\partial x^{*}}\phi_{j}~\mathrm{d}x$. The integrals are computed with a standard Gaussian quadrature rule of three points per finite element. In addition, a classical fixed-point root-finder is utilized to solve for $X_{c}$ in equation (2.13) for a given input Hedstrom number, thus allowing us to compute the yield stress and the nondimensional constants $\lambda_{1}$ and $\lambda_{2}$.
@@
The assembly of the matrices and the solver is handled with the in-house software MAD [64] (used already in several CFD works with finite elements, e.g. [65, 66, 46, 67, 68]), for finite elements and data assimilation, which supports CPU parallelization for optimal speed-up thanks to the underlying linear algebra library PETSc[69]. For this study, we use an AMD Epyc processor with 48 physical cores. We use the MUltifrontal Massively Parallel sparse direct Solver (MUMPs, [70, 71]) to compute the solutions of (2.20), .
@@
The reader may note the absence of stabilization techniques, as it is the gold standard for the discretization of the Navier-Stokes equations. In particular, we remark on the importance of the mass storage component $(1/\Delta t^{*})M$ in the second line of (2.20) as a regularizer that prevents the system of equations from getting close to singular. One should consider nonetheless that getting close to a divergence-free scenario may distort the condition number of the matrix in (2.20). On the other hand, it would be easy to counter a divergence-free situation. There are several stabilization techniques that would take this situation under control, such as the Brezzi-Pitkäranta approach [72], the streamline upwind Petrov-Galerkin one [73], or variational multi-scale methods [74]. Additionally—we can speculate that—since we are not operating in regimes where the convective term is dominant, upwind stabilization strategies are not necessary. The later will be explored in a forthcoming scientific study to test out unstable regimes ensuring numerical resolvability. In addition, the reader may notice that we follow a monolithic scheme for the problem unknowns. Other decoupling strategies for velocity and pressure for this application may be studied in forthcoming research [75].
@@ 3 Results
This section is devoted to study the numerical solutions of the transient slurry equations with convective effects. We first set up, in section 3.1 the necessary mesh and time step for the finite element scheme we employ so that we reach non-oscillatory or not excessively regularized solutions, as a convergence study. Next, in section 3.2, we inspect the numerical solution for a realistic parameter set that emulates copper-water mixture. We then do a parametric study of the solutions including the effect of the fluid yield stress (section 3.3) and the impact of the valve closure time (section 3.4).
@@ 3.1 Convergence study
Let us begin by testing the convergence of the solver in both space and time discretization. For this purpose, we solve the governing equations using the set of parameters shown in Tables 1, 2, and 3, but with $\tau_{y}=0$, as it poses the most challenging numerical scenario. This choice corresponds to sharp time derivatives in the solution and the least dissipative case. Without sufficient refinement, such sharp gradients may lead to numerical oscillations that would otherwise be damped in more viscous regimes. This test is important to assess whether typical stabilization techniques, such as streamline-upwind schemes or modern variational multiscale methods [74], are necessary.
Physically, the most unfavorable scenario concerning the valve closure model, occurs when the closure time is shorter than the round-trip wave travel time along the pipe, i.e., $T_{c}\leq\frac{2L}{c}$. As will be shown in Section 3.4, the limiting case $T_{c}=0$ is numerically challenging, as simulations become prone to oscillations, due to the presence of extremely sharp wave fronts.
Thus, to analyze convergence we set a sudden closure for the valve, meaning that in the boundary conditions of (2.1) we choose $g(t)=0$ at $x=L$. To study the mesh convergence, we narrow our attention to the pressure at a control point located at the valve. Consider a preliminary time step of $\sim 10^{-7}$ s and 4 meshes, with characteristic element sizes of 111 m, 0.20.20.2 m, 0.10.10.1 m, and 0.010.010.01 m. We observe, in Figures 3.a and 3.b, a mesh independent solution up to $\Delta x=0.2$ m, so we choose this element size for all the forthcoming simulations. The critical jump is that of the wave onset, at the beginning of the simulation, so we zoom in the phenomena in Figure 3. For every simulation, we check how the reservoir pressure is recovered after the flow stops, as it is physically expected for the final hydrostatic condition.
(a) Space convergence test
(b) Zoom-in and numerical oscillations
Figure 3: Mesh dependence test. As expected, the mesh refinement is critical after the instantaneous valve closure. This sets a threshold for the element size at $\Delta x=0.2$ m ($L=100m$, with 1001 degrees of freedom). The reader should notice that pressure magnitudes are relative to atmospheric values
We proceed analyzing the time convergence using $\Delta x=0.2$ m. Figure 4.a shows that $\Delta t\geq 2\times 10^{-3}$ s leads to excessively smooth solutions. In addition, we analyze the frequency content of the pressure signal at the valve ($x=L$), as shown in Figure 4.b. We applied a Fast Fourier Transform (FFT) to a $0.48\ \mathrm{s}$ segment of the simulated time history—the duration required for the wave to travel to the reservoir and back. This time‐step selection is adequate: it captures the dominant low‐frequency component near $2.1\ \mathrm{Hz}$ and the subsequent peaks corresponding to the characteristic wave frequency $f=\frac{c}{2L}$.
Due to the fact that we use a constant bulk modulus and omit cavitation, the simulations predict negative pressures well below the fluid’s vapour pressure (as depicted in Figures 3 and 4). In practice, cavitation would prevent such large tensile stresses.
We conclude that $\Delta t=2\times 10^{-5}$ s suffices to recover a coherent and physical solution, therefore we use it for all the simulations to come. In addition, we remark the absence of numerical viscosity achieved at this refinement level, which is a common issue in certain discretization schemes. It worth to mention that one could achieve equally good results with coarser meshes and time steps if we stabilize our finite element solver with classical methods such as the streamline-upwind Petrov-Galerkin approach [76], or modern methods such as the variational multi-scale strategy [77]. This is left as a future work for the authors.
@@
[…]
Concerning the solver performance—which in this test case involves the inversion of 50000 systems of equations with 1001 degrees of freedom (one each time step)—it takes a total simulation time of 66.34 seconds, using a single core of a 48 cores AMD Epyc processor.
@@ 3.2 Simulation of a copper slurry
In this section we will analyze the evolution of a hydraulic transient in a copper-water mixture pipeline system.
Table 1: Physical and mechanical parameters for a copper slurry. Source: [78]
@@
Table 1 shows the physical and mechanical properties for the slurry flow considered in this study, Table 2 the mixture density and bulk modulus and the expected theoretical values for the shockwave speed estimated according to the relations given by Eq.2.9 - Eq.2.11 for increasing values of $C_{V}$. We conduct our simulations using $C_{V}=30\%$. Table 3 display the pipeline properties. The slurry physical properties lead to $\text{He}=1.02\cdot 10^{6}$ [17, 78] and $u_{c}=$ 1.34 $\cdots$-1, then $\text{Re}_{c}=15,390$. For this numerical example the arbitrary value $U_{o}=$ 2.72 m/s was selected as the initial flow velocity. This choice ensures that the initial flow Reynolds number is $\text{Re}=31,257>\text{Re}_{c}=15,390$, placing the system in the turbulent fully-suspended regime.
Table 2: Schockwave speeds ($c$) for a copper tailing according to different authors. In the last column, overpressure values ($\Delta{P}$) were included according to Joukowski formula $\Delta P=\rho_{m}cU_{o}$
@@
Table 3: Parameter setting for the pipeline. The steel pipeline is class DN100 Sch.40s [79]
@@
In agreement with collected experience in engineering projects, results provided by Eqs. (2.9)-(2.10) are extremely large and thus unrealistic, and may need experimental verification in the context of copper slurry transients. The wave speed values calculated from Eq. (2.11) can be considered most representative of a real transient event, in agreement with Kodura et al. [7, 8, 9]. Even more, it is expected that real values of the initial pressure surge are lower than theoretical estimations inspired on Joukowski’s pressure scale, because of the lower shockwave speed (see e.g., [7, 9]). Thus, all our numerical simulations consider the wave speed computed from Eq. (2.11)
With these elements, the Joukowsky’s pressure rise can be easily estimated from Eq.2.7 as $\Delta{P}=\rho_{m}cU_{o}$. The last column of Table 2 shows some estimations of $\Delta P$ by considering the reference velocity $U_{o}$. By one hand, the maximum pressure values along the pipeline $p_{max}(x)$ obtained from numerical simulations can be compared to the theoretical maximum pressure $\max_{0\leq x\leq L}(p(x,t=0)+\Delta P)$, where $p(x,t=0)$ is the steady-state pressure profile of the pipeline. On the other hand, $p_{max}(x)$ can also be compared to $P_{A}$, where $P_{A}$ is the allowable pressure. As expected, a risky situation for pipeline safety occurs when $P_{max}\geq P_{A}$. This last parameter is provided by the pipeline factory and it depends on the wall material and schedule (see Table 3). In addition, the minimum pressures in the system, $P_{min}(x)$, can also cause problems in the pipeline. In this case, the risk situation will occur when $P_{min}(x)<P_{v}$, where $P_{v}$ is the vacuum pressure that induces the liquid-gas phase separation in the mixture. A typical value for this parameter is $P_{v}=-51.32$ Pa (using a gauge pressure scale).
@@
In addition, we show the temporal evolution of the numerical solutions for control points located at the tank ($x=0$), the pipe half length ($x=L/2$), and the valve ($x=L$), at Figures 6 and 7. We observe a rather slow friction dissipation and we check that the maximal numerical pressure reach a value of 13.8613.8613.86 MPa, which turns out to be surprisingly close to the factory allowable pressure $P_{A}=17.8$ MPa, but still ensuring a safe hammer wave propagation. The theoretical pressure rise discussed above, with more restricted assumptions, is lower than the numerical one, namely 111111 MPa, yet in accordance with the obtained value.